# Unconventional early-time relaxation in the Rydberg chain

### v1

@@ Abstract
We show that unconventional relaxation dynamics of special initial states in one-dimensional arrays of Rydberg atoms produce non-generic decay of the initial-state survival probability at early times. Using the PXP hamiltonian as a minimal model of the Rydberg blockade, we prove that the early-time survival probability for states exhibiting quantum many-body scarring decays at a characteristic rate, whose finite-size scaling is determined solely by scars. We numerically investigate the effects of both revival-enhancing and ergodicity-restoring perturbations and find results consistent with the limiting cases of integrable and ergodic dynamics, respectively. Since the survival probability is easily accessible experimentally at early times, our findings enable us to probe the presence of scars at time scales much shorter than that of thermalization.
Since its discovery in a Rydberg atom array [1], weak ergodicity breaking in quantum systems with constrained dynamics is under active investigation [2, 3, 4, 5, 6]. This phenomenon has been associated with quantum many-body scars (QMBSs), special non-thermal mid-spectrum eigenstates. One way to detect the presence of QMBSs is to quench a fine-tuned initial state with high overlap with QMBS eigenstates and look for oscillations of observables that persist for times much longer than the expected thermalization time of the system. Ascertaining the presence of QMBSs via observation of long-lived oscillations can be challenging due to the short coherence windows of existing quantum devices. Uncovering early-time signatures of QMBSs would therefore render weak ergodicity breaking and associated effects more amenable to noisy quantum simulation.
In this Letter, we trace the presence of QMBSs in the survival probability (SP) of initial states leading to scarring. The SP  [^1] Also called fidelity, nondecay probability, or return probability. is defined as the overlap between an initial state $|\psi_{0}\rangle$ and the evolved state $|\psi_{t}\rangle$ at time $t$:
@@
The rate $\sigma^{2}$ is characteristic of the underlying dynamics and the initial state, and provides global information about the stability of quantum dynamics in isolated interacting quantum systems [9, 10]. For evolution under a hamiltonian $H$, $\sigma^{2}$ is determined by the spread of the initial state over the energy eigenbasis $\{|E_{\alpha}\rangle\}$ of $H$ as
@@
where .$\langle.\rangle_{\psi_{0}}$ denotes expectation with respect to $|\psi_{0}\rangle$, $E_{0}=\langle H\rangle_{\psi_{0}}$ is the energy of the initial state, and $C_{\alpha}^{0}=\langle\psi_{0}|E_{\alpha}\rangle$. In other words, the decay rate $\sigma^{2}$ is the variance of the overlap distribution $P_{\alpha}^{0}=\big{|}C_{\alpha}^{0}\big{|}^{2}$ of the initial state over energy eigenstates. The SP is then
$\mathcal{S}_{\text{P}}(t)=\bigg{|}\,\sum_{\alpha}\big{|}C_{\alpha}^{0}\big{|}^{2}e^{-itE_{\alpha}}\,\bigg{|}^{2}\,.$ (5)
@@
In the thermodynamic limit, the discrete probability distribution $P_{\alpha}^{0}$ becomes $\mathcal{P}{\scriptscriptstyle\text{LDOS}}(E)=\sum_{\alpha}|C_{\alpha}^{(0)}|^{2}\;\delta(E-E_{\alpha})$, known as the local density of states (LDOS). For simplicity, we use the term LDOS for both the discrete and the continuous case.
@@
In generic quantum many-body systems with 2-local interactions, initial states with energy close to the middle of the spectrum exhibit a Gaussian LDOS. It has been shown that for both chaotic and strongly perturbed integrable systems, the decay rate scales with system size $L$ as $\sigma^{2}\propto L$ [9, 10, 11].
@@
In contrast, below we show that QMBS systems may exhibit a non-generic SP decay rate that is fully determined by the QMBS component of the initial-state LDOS alone. The intuition is that the LDOS of initial states leading to scarred dynamics is dominated by a comb of QMBS contributions (see, e.g., Fig. 1(b)), which fix $\sigma^{2}$. Assuming a semi-phenomenological model of a scarred LDOS, we derive general conditions under which the presence of QMBSs in the energy spectrum can be probed by measuring the early-time decay of the SP. The SP thus provides a sought-after early-time signature of QMBSs accessible in current rapidly-decohering quantum simulators.
@@
We demonstrate this reasoning in the PXP hamiltonian, an effective model for a chain of $L$ neutral atoms (qubits) subjected to Rydberg blockade [1], defined as
$H_{\text{PXP}}=\sum_{i=1}^{L}P_{i-1}X_{i}P_{i+1}\,,$ (6)
@@
where $X_{i}$, $Z_{i}$ are Pauli matrices associated to the $i$-th qubit, and $P_{i}=(1-Z_{i})/2=\outerproduct{1}{1}$ |. It is also useful to define $Q_{i}=1-P_{i}=\outerproduct{0}{0}$ |. In what follows, we consider only periodic boundary conditions.
@@
Figure 1: LDOS of the |$\ket{\mathbb{Z}_{2}}$ state over the energy eigenbasis of PXP model (1(a)) with revival-enhancing deformation ($g_{\text{perf}}=0.048$), (1(b)) without deformation, and (1(c)) with ergodicity-restoring deformation ($g_{\text{NNN}}=1$). The black dashed line is a Gaussian fit for the LDOS. All plots are obtained by Lanczos exact diagonalization for $L=28$ qubits with 400 iterations (histogrammed with 130 bins).
@@
A quench of the state |$\ket{\mathbb{Z}_{2}}=\ket{1010\dots}$ with the PXP hamiltonian leads to long-lived revivals of the state. This behavior can be traced back to the significant overlap of the |$\ket{\mathbb{Z}_{2}}$ state with $k=L-1$ QMBS eigenstates, leading to $L-1$ narrow peaks aligned with QMBS energies, approximately equally separated by a size-independent energy gap [12, 13] — see Fig. 1(b). The revival period is inversely proportional to this gap. The resulting LDOS differs significantly from the usual Gaussian distribution observed in generic evolution (see, e.g., Fig. 1(c)).
@@
The variance of the full LDOS can be analytically calculated from Eq. (3), and is given by $\sigma_{\text{PXP}|\mathbb{Z}_{2}}^{2}=L/2$, since $\langle H_{\mathrm{PXP}}\rangle_{\mathbb{Z}_{2}}=0$ and only $L/2$ terms contribute to the expectation $\langle H_{\mathrm{PXP}}^{2}\rangle_{\mathbb{Z}_{2}}$. This non-generic $L/2$ scaling distinguishes the scarred LDOS from the generic Gaussian case [9, 10, 11]. In contrast, an analogous calculation for the rapidly thermalizing initial state |$\ket{\mathbb{Z}_{1}}=\ket{1}^{\otimes L}$ yields the generic behaviour $\sigma_{\text{PXP}|\mathbb{Z}_{1}}^{2}=L$.
@@
We now show that the non-generic scaling of $\sigma^{2}$ is determined exclusively by QMBSs in the PXP model, thus signalling their presence. An excellent approximation of QMBS eigenstates can be obtained using the so-called forward-scattering approximation (FSA) [13, 14]. The FSA allows us to numerically isolate QMBS contributions to the LDOS. The FSA calculation yields a variance $\sigma_{\text{PXP}|\mathbb{Z}_{2}}^{2\,\text{(FSA)}}=L/2$ — see Fig. 2.
@@
To further reinforce the conclusion that scars alone, if present, determine the early-time SP in the PXP model, we study the effect of perturbations $\delta H$ that either enhance revivals or restore ergodicity. In what follows, we only consider perturbations and initial states for which $\langle\delta H\rangle_{\psi_{0}}=0$. Thus, upon introducing such a perturbation, the LDOS variance for the total hamiltonian $H=H_{\text{PXP}}+\delta H$ becomes
@@
Below we evaluate this expression for |$\ket{\psi_{0}}=\ket{\mathbb{Z}_{2}},\ket{\mathbb{Z}_{1}}$ and two different perturbations.
We first introduce a term known to enhance the fidelity of the revivals [15, 16]. This reads
$H_{\text{perf}}=g_{\text{perf}}\sum_{j}P_{j-1}X_{j}P_{j+1}(Z_{j+2}+Z_{j-2})\,,$ (8)
with $g_{\text{perf}}\geq 0$. Its effect is to increase the rate of flips on qubit $j$ when a qubit is in $|0\rangle$ two sites away. Nearly perfect revivals are obtained for $g_{\text{perf}}\approx 0.048$.
The |$\ket{\mathbb{Z}_{2}}$ LDOS with respect to the total hamiltonian $H$ with $\delta H=H_{\text{perf}}$ remains highly peaked within the range $g_{\text{perf}}\in[0,0.048]$, with the number of peaks staying constant. For the |$\ket{\mathbb{Z}_{2}}$ initial state, the terms in Eq. (7) are evaluated as $\langle H_{\text{PXP}}H_{\text{perf}}\rangle_{\mathbb{Z}_{2}}=\langle H_{\text{perf}}H_{\text{PXP}}\rangle_{\mathbb{Z}_{2}}=g_{\text{perf}}L$ and $\langle H_{\text{perf}}^{2}\rangle_{\mathbb{Z}_{2}}=2g_{\text{perf}}^{2}L$. The LDOS variance is therefore $\sigma_{\text{perf}|\mathbb{Z}_{2}}^{2}=(2g_{\text{perf}}^{2}+2g_{\text{perf}}+\frac{1}{2})L$. The scar variance extracted from FSA numerics agrees perfectly with this scaling, as shown in Fig. 2. In contrast, the LDOS variance of the |$\ket{\mathbb{Z}_{1}}$ state remains Gaussian without dominant overlap with scars. The scaling of its variance is distinct from that of |$\ket{\mathbb{Z}_{2}}$. We namely find $\sigma_{\text{perf}|\mathbb{Z}_{1}}^{2}=(4g_{\text{perf}}^{2}-4g_{\text{perf}}+1)L$. In the range $g_{\text{perf}}\ll 1$, where revival enhancement is observed for the |$\ket{\mathbb{Z}_{2}}$ initial state, we notice that $H_{\text{perf}}$ has the opposite effect on the LDOS variance of the |$\ket{\mathbb{Z}_{2}}$ and |$\ket{\mathbb{Z}_{1}}$ initial states.
@@
As can be seen in Fig. 1(c), $H_{\text{NNN}}$ suppresses the peaks in the |$\ket{\mathbb{Z}_{2}}$ LDOS and a generic Gaussian LDOS is observed for large enough $g_{\text{NNN}}$. We note that restoration of ergodicity means that the FSA is no longer valid. The exact expression for the variance scaling is $\sigma_{\text{NNN}|\mathbb{Z}_{2}}^{2}=(\frac{1}{2}g_{\text{NNN}}^{2}+\frac{1}{2})L$. The |$\ket{\mathbb{Z}_{1}}$ LDOS behaves similarly in this case, with scaling $\sigma_{\text{NNN}|\mathbb{Z}_{1}}^{2}=(g_{\text{NNN}}^{2}+1)L$. This is to be contrasted with the opposite trends obtained above for |$\ket{\mathbb{Z}_{2}}$ and |$\ket{\mathbb{Z}_{1}}$ initial states under the scar-enhancing deformation.
@@
Building on the insights collected so far, we devise a general semi-phenomenological model of a scarred LDOS for initial states having overlap with both QMBS eigenstates and otherwise thermal eigenstates. It follows from the early idea that the energy spectrum of scarred models can be divided in a subspace of QMBS eigenstates and a subspace of thermal eigenstates [17]. The relevant initial states to probe the presence of QMBS will generally be those having an overlap over both sectors, but whose overlap with the QMBS sector is significantly higher than with the thermal one.
@@
To motivate our scarred LDOS model, we consider the sum of the bare PXP hamiltonian, hosting approximate QMBS, and any perturbation $\delta\tilde{H}$ for which $\langle\delta\tilde{H}\rangle_{\psi_{0}}=\langle H_{\text{PXP}}\delta\tilde{H}\rangle_{\psi_{0}}=\langle\delta\tilde{H}H_{\text{PXP}}\rangle_{\psi_{0}}=0$ for the initial state(s) of interest, such as the $H_{\text{NNN}}$ introduced above. Due to the linearity of the expectation, this is the case whenever there are no terms in $\delta\tilde{H}$ that can undo the effect of the terms of $H_{\text{PXP}}$ on the initial state. Since each term of the latter contains an odd number of $X$s (one here), the local terms of the former must contain an odd number of them as well to cancel their effect. Therefore, for these perturbations, the variance reduces to the sum of the variances of PXP and deformation terms: $\sigma^{2}=\langle H_{\text{PXP}}^{2}\rangle+\langle\delta\tilde{H}^{2}\rangle$. Specifically, the effect of the ergodicity-restoring deformation $H_{\text{NNN}}$, whose terms are parity-preserving, can be seen as an independent contribution to the LDOS variance.
@@
With the above reasoning in mind, we can abstract away all microscopic model details and focus on LDOS of the form
$\mathcal{P}_{\scriptscriptstyle\text{LDOS}}(E)=w_{1}\,\mathcal{P}_{\scriptscriptstyle\text{SC}}(E)+w_{2}\,\mathcal{P}_{\scriptscriptstyle\text{TH}}(E)$ (10)
@@
with the constraint $w_{1}+w_{2}=1$ to ensure normalization. The distribution $\mathcal{P}_{\scriptscriptstyle\text{SC}}(E)$ describes (approximate) scars and $\mathcal{P}_{\scriptscriptstyle\text{TH}}(E)$ represents the background of thermal states, both normalized to 1. Motivated by our results for the deformed PXP model discussed above, we assume that both distributions have the same mean, so that $\mu_{\text{LDOS}}=\mu_{\text{SC}}=\mu_{\text{TH}}$, and, further, that $\sigma_{\text{SC}}^{2}\neq\sigma_{\text{TH}}^{2}$, where $\sigma_{\text{SC}}^{2}$ and $\sigma_{\text{TH}}^{2}$ are the variances of the distributions of each sector. The total LDOS variance is then
$\sigma_{\text{LDOS}}^{2}=w_{1}\sigma_{\text{SC}}^{2}+w_{2}\sigma_{\text{TH}}^{2}\,.$ (11)
In this setting, $\sigma_{\text{SC}}^{2}$ fully determines the observed $\sigma_{\text{LDOS}}^{2}$ if $w_{1}\sigma_{\text{SC}}^{2}\gg w_{2}\sigma_{\text{TH}}^{2}$, i.e., if
$\frac{w_{1}}{w_{2}}\gg\frac{\sigma_{\text{TH}}^{2}}{\sigma_{\text{SC}}^{2}}\,,$ (12)

### v2

@@ Abstract
We show that unconventional relaxation dynamics of special initial states in one-dimensional arrays of Rydberg atoms produce non-generic decay of the initial-state survival probability (SP) at very early times. Using the PXP hamiltonian as a minimal model of the Rydberg blockade, we prove that the early-time SP for states exhibiting quantum many-body scarring (QMBS) decays at a characteristic rate, whose finite-size scaling is determined solely by scars. We numerically investigate the effects of both revival-enhancing and ergodicity-restoring deformations of the PXP hamiltonian and find results consistent with the limiting cases of integrable and ergodic dynamics, respectively. We moreover argue that such unconventional early relaxation of scarred initial states is characteristic of a whole class of QMBS models. Since the SP can be easily measured experimentally, our findings enable us to probe the presence of scars at time scales much shorter than that of thermalization.
Since its discovery in a Rydberg atom array [1], weak ergodicity breaking in quantum systems with constrained dynamics is under active investigation [2, 3, 4, 5, 6]. This phenomenon has been associated with quantum many-body scars (QMBSs), special non-thermal mid-spectrum eigenstates. They have been shown to be present in varying number in a wide range of models, going from $O(1)$ to a number polynomial and even exponential in the system size.
So far, experimental observations of QMBS have been achieved only for models with multiple scars. In this case their presence can be detected by quenching a fine-tuned initial state having high overlap with the QMBS and look for oscillations that persist for long times, much longer than the expected thermalization time of the system. However, this is challenging in quantum simulators with short coherence windows. Moreover, certain scarred dynamics do not lead to coherent revivals, an obvious example being models with a single scar. Uncovering signatures of QMBSs accessible at much earlier times and for a broader range of models would therefore render weak ergodicity breaking and associated effects more amenable to noisy quantum simulation.
In this Letter, we trace the presence of QMBSs in the early-time decay of survival probability (SP) of initial states leading to scarring. The SP  [^1] Also called revival fidelity, nondecay probability, or return probability. is defined as the overlap between an initial state $|\psi_{0}\rangle$ and the evolved state $|\psi_{t}\rangle$ at time $t$:
@@
The decay rate $\sigma^{2}$ is characteristic of the underlying dynamics and the initial state, and provides global information about the stability of quantum dynamics in isolated interacting quantum systems [8, 9]. For evolution under a hamiltonian $H$, $\sigma^{2}$ is determined by the spread of the initial state over the energy eigenbasis $\{|E_{\alpha}\rangle\}$ of $H$ as
@@
where .$\langle.\rangle_{\psi_{0}}$ denotes expectation with respect to $|\psi_{0}\rangle$, $E_{0}=\langle H\rangle_{\psi_{0}}$ is the energy of the initial state, and $C_{\alpha}^{0}=\langle\psi_{0}|E_{\alpha}\rangle$. In other words, the decay rate $\sigma^{2}$ is the variance of the overlap distribution $P_{\alpha}^{0}=\big{|}C_{\alpha}^{0}\big{|}^{2}$ of the initial state over energy eigenstates. In the thermodynamic limit, the discrete probability distribution $P_{\alpha}^{0}$ becomes $\mathcal{P}{\scriptscriptstyle\text{LDOS}}(E)=\sum_{\alpha}|C_{\alpha}^{(0)}|^{2}\;\delta(E-E_{\alpha})$, known as the local density of states (LDOS). For simplicity, we use the term LDOS for both the discrete and the continuous case.
@@
In generic quantum many-body systems with local interactions, initial states with energy close to the middle of the spectrum exhibit a Gaussian LDOS. It has been shown that their variance, and thus the SP decay rate, generically scales with system size $L$ as
$\sigma_{\text{gen.}}^{2}\propto J^{2}L$ (5)
where $J$ is the main energy scale of the system [8, 9, 10].
@@
In contrast, we argue below that the LDOS of scarred initial states generally diverges substancially from this Gaussian form, causing a non-generic SP decay that can be seen at times significantly shorter than those usually considered to probe scarring effects. Beyond perfect scars, we show it holds in a paradigmatic model of approximate scars, the PXP model, which is also of experimental relevance due to its proximity with Rydberg-atom experiments. In both cases, the intuition is that the LDOS of initial states leading to scarred dynamics is dominated by a comb of QMBS contributions (see, e.g., Fig. 1(b)), which fix $\sigma^{2}$. Assuming a semi-phenomenological model of a scarred LDOS, we derive general conditions under which the presence of QMBSs in the energy spectrum can be probed by measuring the early-time decay of the SP. Finally, we broaden the scope of our study by also distinguishing the cases of a single scar and of a few scars. The SP initial decay thus provides a sought-after early-time signature of QMBSs accessible in current rapidly-decohering quantum simulators.
@@
The PXP hamiltonian is an effective model for a chain of $L$ neutral atoms (qubits) subjected to Rydberg blockade [1] (see SM for a derivation), and is defined as
$H_{\text{PXP}}=J\sum_{i=1}^{L}P_{i-1}X_{i}P_{i+1}\,,$ (6)
@@
where $J=\frac{\Omega}{2}$, $\Omega$ is the Rabi frequency, $X_{i}$, $Z_{i}$ are Pauli matrices associated to the $i$-th qubit, and projectors $P_{i}=(1-Z_{i})/2=\outerproduct{1}{1}$ | enforce the Rydberg blockade constraint. We consider periodic boundary conditions.
@@
Figure 1: LDOS of the |$\ket{\mathbb{Z}_{2}}$ state over the energy eigenbasis of PXP model with $J=1$ (1(a)) with revival-enhancing deformation ($g_{\text{PXPZ}}=0.048$), (1(b)) without deformation, and (1(c)) with ergodicity-restoring deformation ($g_{\text{NNN}}=1$). The black dashed line is a Gaussian fit for the LDOS. All plots are obtained by Lanczos exact diagonalization for $L=28$ qubits with 400 iterations (histogrammed with 130 bins).
@@
A quench of the state |$\ket{\mathbb{Z}_{2}}=\ket{1010\dots}$ with the PXP hamiltonian leads to long-lived revivals of the initial state. This behavior can be traced back to the significant but approximate overlap of the |$\ket{\mathbb{Z}_{2}}$ state with $k=L-1$ QMBS eigenstates. The |$\ket{\mathbb{Z}_{2}}$ LDOS takes the form of $L-1$ narrow peaks aligned with QMBS energies, approximately equally separated by a size-independent energy gap [11, 12] — see Fig. 1(b). The revival period is inversely proportional to this gap. The resulting LDOS differs significantly from the usual Gaussian distribution observed in generic evolution (see, e.g., Fig. 1(c)).
@@
The variance of the full LDOS can be analytically calculated from Eq. (3), and is given by $\sigma_{\text{PXP}|\mathbb{Z}_{2}}^{2}=J^{2}\frac{L}{2}$, since $\langle H_{\mathrm{PXP}}\rangle_{\mathbb{Z}_{2}}=0$ and only $L/2$ terms contribute to the expectation $\langle H_{\mathrm{PXP}}^{2}\rangle_{\mathbb{Z}_{2}}$. This non-generic $L/2$ scaling distinguishes the scarred LDOS from the generic Gaussian case [8, 9, 10]. In contrast, an analogous calculation for the rapidly thermalizing initial state |$\ket{\mathbb{Z}_{1}}=\ket{1}^{\otimes L}$ yields the generic behaviour $\sigma_{\text{PXP}|\mathbb{Z}_{1}}^{2}=J^{2}L$.
[…]
@@
We now show that the non-generic scaling of $\sigma^{2}$ is determined exclusively by QMBSs in the PXP model, thus signalling their presence. An excellent approximation of QMBS eigenstates can be obtained using the so-called forward-scattering approximation (FSA) [12, 13]. The FSA allows us to numerically isolate QMBS contributions to the LDOS. The FSA calculation yields a variance $\sigma_{\text{PXP}|\mathbb{Z}_{2}}^{2\,\text{(FSA)}}=L/2$ — see Fig. 2 (we take $J=1$ for all the numerics).
@@
To further reinforce the conclusion that scars alone, if present, determine the early-time SP in the PXP model, we study the effect of deformations $\delta H$ that either enhance revivals or restore ergodicity. In what follows, we only consider deformations and initial states for which $\langle\delta H\rangle_{\psi_{0}}=0$. Thus, upon introducing such a deformation, the LDOS variance for the total hamiltonian $H=H_{\text{PXP}}+\delta H$ becomes
@@
Below we evaluate this expression for |$\ket{\psi_{0}}=\ket{\mathbb{Z}_{2}},\ket{\mathbb{Z}_{1}}$ and two different deformations. The effect of these perturbations on the |$\ket{\mathbb{Z}_{2}}$ LDOS are shown in Fig. 1(a),1(c).
We first introduce a term known to enhance the fidelity of the revivals in the perturbative regime [14, 15]. This reads
$H_{\text{PXPZ}}=g_{\text{PXPZ}}\sum_{j}P_{j-1}X_{j}P_{j+1}(Z_{j+2}+Z_{j-2})\,,$ (8)
with $g_{\text{PXPZ}}\geq 0$. Its effect is to increase the rate of flips on qubit $j$ when a qubit is in $|0\rangle$ two sites away. Nearly perfect revivals are obtained for $g_{\text{PXPZ}}\approx 0.048$.
The |$\ket{\mathbb{Z}_{2}}$ LDOS with respect to the total hamiltonian $H$ with $\delta H=H_{\text{PXPZ}}$ remains highly peaked within the range $g_{\text{PXPZ}}\in[0,0.048]$, with the number of peaks staying constant. For the |$\ket{\mathbb{Z}_{2}}$ initial state, the terms in Eq. (7) are evaluated as $\langle H_{\text{PXP}}H_{\text{PXPZ}}\rangle_{\mathbb{Z}_{2}}=\langle H_{\text{PXPZ}}H_{\text{PXP}}\rangle_{\mathbb{Z}_{2}}=Jg_{\text{PXPZ}}L$ and $\langle H_{\text{PXPZ}}^{2}\rangle_{\mathbb{Z}_{2}}=2g_{\text{PXPZ}}^{2}L$. The LDOS variance is therefore $\sigma_{\text{PXPZ}|\mathbb{Z}_{2}}^{2}=(4g_{\text{PXPZ}}^{2}+4Jg_{\text{PXPZ}}+J^{2})\frac{L}{2}$. The scar variance extracted from FSA numerics agrees perfectly with this scaling, as shown in Fig. 2. In contrast, the LDOS variance of the |$\ket{\mathbb{Z}_{1}}$ state remains Gaussian without dominant overlap with scars. The scaling of its variance is distinct from that of |$\ket{\mathbb{Z}_{2}}$. We namely find $\sigma_{\text{PXPZ}|\mathbb{Z}_{1}}^{2}=(4g_{\text{PXPZ}}^{2}-4Jg_{\text{PXPZ}}+J^{2})L$. In the range where revival enhancement is observed for the |$\ket{\mathbb{Z}_{2}}$ initial state, $g_{\text{PXPZ}}\ll 1$, we notice that $H_{\text{PXPZ}}$ has an opposite effect on the LDOS variance of the |$\ket{\mathbb{Z}_{2}}$ and |$\ket{\mathbb{Z}_{1}}$ initial states.
@@
As can be seen in Fig. 1(c), $H_{\text{NNN}}$ suppresses the peaks in the |$\ket{\mathbb{Z}_{2}}$ LDOS and a generic Gaussian LDOS is observed for large enough $g_{\text{NNN}}$. We note that restoration of ergodicity means that the FSA is no longer valid, even for small $g_{\text{NNN}}$. The exact expression for the variance scaling is $\sigma_{\text{NNN}|\mathbb{Z}_{2}}^{2}=(g_{\text{NNN}}^{2}+J^{2})\frac{L}{2}$. The |$\ket{\mathbb{Z}_{1}}$ LDOS behaves similarly in this case, with scaling $\sigma_{\text{NNN}|\mathbb{Z}_{1}}^{2}=(g_{\text{NNN}}^{2}+J^{2})L$. This is to be contrasted with the opposite trends obtained above for |$\ket{\mathbb{Z}_{2}}$ and |$\ket{\mathbb{Z}_{1}}$ initial states under the scar-enhancing deformation.
@@
This intuition can be confirmed analytically for a class of models typically hosting equidistant QMBS eigenstates, the so-called spectrum-generating algebra (SGA) class [16]. This framework leads to hamiltonians with the form, $H=J\ H^{\prime}+h\ Q^{z}$, where there are distinct energy scales $h$ for the scar sector forming a basis of the operator $Q^{z}$ and $J$ for the thermal sector associated with $H^{\prime}$. In these models, relevant initial states have been found that are superpositions of the QMBS only [17, 18, 19, 20]. They lead to scarring dynamics that is fully disconnected from the thermal sector of the energy spectrum. The early-time SP decay rate of these initial states is shown to depend exclusively on QMBS features. A simple example is provided by the scarred spin-1 XY model [19, 21], defined as $H=J\ H_{\text{XY}}+h\ (2J^{z})=J\ \sum_{i=1}^{L}(S_{i}^{x}S_{i+1}^{x}+S_{i}^{y}S_{i+1}^{y})+h\ \sum_{i=1}^{L}S_{i}^{z}$, and hosting $L+1$ equally spaced scar eigenstates. Considering a simple initial state that is a superposition of QMBS leads to an early-time SP decay rate $\sigma_{\text{XY}|\text{scar}}^{2}=h^{2}L$ controlled only by the energy scale $h$ of the scarred part of the hamiltonian. In contrast, the decay rate of a generic initial product state leads to a decay rate controlled by the energy scale of the nonintegrable part of the hamiltonian, $\sigma_{\text{XY}|\text{gen.}}^{2}=J^{2}L$. Here again the early decay time scale is much shorter than the time scale of the first revival and the separation increases as $\sqrt{L}$. Note that this spin-1 XY model with QMBS can be generalized to non-translation invariant model, and to higher spatial dimensions (see the SM for details).
@@
Building on the insights collected so far, we shed light on the conditions under which the presence of QMBSs can be probed by measuring the early-time decay of the SP using a general model of a scarred LDOS. This semi-phenomenological model follows from the early idea that the energy spectrum of scarred models can be divided in a subspace of QMBS eigenstates and a subspace of thermal eigenstates [16], $H\approx\mathcal{H}_{\text{scar}}\oplus\mathcal{H}_{\text{thermal}}$. The relevant initial states to probe the presence of QMBS will generally be those having an overlap over both sectors, but whose overlap with the QMBS sector is significantly higher than with the thermal one.
@@
To motivate our scarred LDOS model, we consider the sum of the bare PXP hamiltonian, hosting approximate QMBSs, and any deformation $\delta\tilde{H}$ for which $\langle\delta\tilde{H}\rangle_{\psi_{0}}=\langle H_{\text{PXP}}\delta\tilde{H}\rangle_{\psi_{0}}=\langle\delta\tilde{H}H_{\text{PXP}}\rangle_{\psi_{0}}=0$ for the initial state(s) of interest, such as the $H_{\text{NNN}}$ deformation introduced above. Due to the linearity of the expectation, this is the case whenever there are no terms in $\delta\tilde{H}$ that can undo the effect of the terms of $H_{\text{PXP}}$ on the initial state. Since each term of the latter contains an odd number of $X$s (one here), the local terms of the former must contain an odd number of them as well to cancel their effect. Therefore, for these deformations, the variance reduces to the sum of the variances of PXP and deformation terms: $\sigma^{2}=\langle H_{\text{PXP}}^{2}\rangle+\langle\delta\tilde{H}^{2}\rangle$. Specifically, the effect of the ergodicity-restoring deformation $H_{\text{NNN}}$, whose terms are parity-preserving, can be seen as an independent contribution to the LDOS variance.
@@
With the above reasoning in mind, we can abstract away all microscopic model details and focus on a simple model of scarred LDOS of the form
$\mathcal{P}_{\text{LDOS}}(E)=w_{\text{scar}}\,\mathcal{P}_{\text{scar}}(E)+w_{\text{th.}}\,\mathcal{P}_{\text{th.}}(E)$ (10)
@@
with the constraint $w_{\text{scar}}+w_{\text{th.}}=1$ to ensure normalization. The distribution $\mathcal{P}_{\text{scar}}$ describes (approximate) scars and $\mathcal{P}_{\text{th.}}$ represents the background of thermal states, both normalized to 1.
Motivated by our results for the deformed PXP model discussed above, we assume that both distributions have the same mean, so that $\mu_{\text{LDOS}}=\mu_{\text{scar}}=\mu_{\text{th.}}$. However, we consider the general case for which $\sigma_{\text{scar}}^{2}\neq\sigma_{\text{th.}}^{2}$, where $\sigma_{\text{scar}}^{2}$ and $\sigma_{\text{th.}}^{2}$ are the variances of the distributions of each sector. The total LDOS variance is then $\sigma_{\text{LDOS}}^{2}=w_{\text{scar}}\sigma_{\text{scar}}^{2}+w_{\text{th.}}\sigma_{\text{th.}}^{2}$. In this general setting, one finds that $\sigma_{\text{scar}}^{2}$ fully determines the observed $\sigma_{\text{LDOS}}^{2}$, i.e., $\sigma_{\text{LDOS}}^{2}\simeq\sigma_{\text{scar}}^{2}$ if
$\frac{w_{\text{scar}}}{w_{\text{th.}}}\gg\frac{\sigma_{\text{th.}}^{2}}{\sigma_{\text{scar}}^{2}},$ (11)